A market consists of 5 firms, each of which has 20% market share. Then the Herfindahl
index (HHI) is equal to:
A. 0.1. B. 0.15. C. 0.2. D. 0.25. E. 0.3.
F. none of the above.

PLEASE include explanation
Answer #1

C. 0.2

Explanation: HHI is the sum total of the squares of the market shares of the firms = (0.2)2 + (0.2)2 + (0.2)2 + (0.2)2​​​​​​​​​​​​​​ + (0.2)2 = 0.04 + 0.04 + 0.04 + 0.04 + 0.04 = 0.20.